Saturday, 22 August

Plenary Lecture

9:30 – 10:30  Auditorium Maximum H 105

Wolfgang Hackbusch, Universität Kiel, Kiel, Germany From Classical Numerical Mathematics to Scientific Computing (I: p.6)

Chair: Willi Jäger, Universität Heidelberg, Heidelberg, Germany

ICM’98 – Program  1